4. Mr. Contreras knows that his margins in his gasoline station is just 2.6%. If the markup amount per liter of diesel is around php 1.23, how much is the cost per liter of diesel?what is the selling price per liter of diesel?





5. A house and lot with original price of php 2,300,000 was sold at discounted price of php 2,150,000. What was the discount rate given to the buyer?



1
Expert's answer
2021-11-26T13:53:17-0500

4) Selling Price "=1.23\div2.6\%=47.31"

 cost per liter of diesel"=47.3-1.23=46.08"

5) Original price "=php 2,300,000"

Selling price "=php 2,150,000"

Discount "= 2,300,000-2,150,000=php 150,000"

The discount rate  is given by the following formula

discount rate "={Discount\over Original price}\times 100\%={php 150,000\over php 2,300,000}\times 100\%=6.5217\%"